# thanks: These authors contributed equally to this workthanks: These authors contributed equally to this work

### v1

@@ B.2 The mixed state case
Using the default settings of the programs implemented by MATLAB, all the global optimization algorithms are executed on a computer with an Intel i9-12900KS Core and a RAM of 128 GB for 24 hours. Meanwhile, an MLP model consisting of a 7-layer network with each layer containing 90 neurons is also trained 10 times with the other configuration remaining the same as that of VGON. The numerical results are shown in Table. 2, where we can clearly see that the performance of VGON exceeds that of all the other methods.
Table 2: Comparisons of VGONs with seven global optimization algorithms and MLP.
@@
MLP 0.0384 3000 iterations
VGON 0.0837 3000 iterations
Appendix C Alleviating the effect of barren plateaux in variational quantum algorithms
@@ C.1 The $Z_{1}Z_{2}$ model
is adopted. Here $\theta_{l}^{i}\in[0,2\pi)$ are variational angles, and all the $L\times N$ such angles combined is denoted as $\bm{\theta}$. The rotations $R_{l}^{i}(\theta_{l}^{i})=\exp(-\frac{\rm{i}}{2}\theta_{l}^{i}G_{l}^{i})$ have random directions given by $G_{l}^{i}\in\{\sigma_{x},\sigma_{y},\sigma_{z}\}$. $U_{ENT}$ is an entangling unitary operation consisting of two-qubit nearest-neighbor controlled-Z (CZ) gates with periodic boundary conditions. $L$ and $N$ correspond to the numbers of the layers and qubits of $U(\bm{\theta})$ respectively. The structure of $U(\bm{\theta})$ for the $Z_{1}Z_{2}$ model is schematically shown in Fig. 7.
Figure 7: Structure of the parameterized quantum circuit for the $Z_{1}Z_{2}$ model. Each dark-blue wireframe represents a layer of the circuit consisting of single-qubit rotations represented by light-blue boxes and an entangling unitary operation $U_{ENT}$ represented as a purple box, where entangling CZ gates are shown by lines. The measurements at the end are used to estimate the energy of the trial state.
Let the numbers of qubits and layers be $20$ and $400$ respectively. To approximate the ground states of the $Z_{1}Z_{2}$ model by VQE, the variable $\bm{\theta}\in\mathbb{R}^{8,000}$ is uniformly initialized from the parameter space, i.e., $\theta_{l}^{i}\in[0,2\pi)$, and then updated iteratively by the Adam optimizer [28] to minimize $h(\bm{\theta})$. After 300 iterations, the obtained energy is -0.0068, which is actually far from the ground energy -1. Particularly, the variance of $\{\partial_{{\theta_{l}^{i}}}h(\bm{\theta})\}$ decreases from $4.4999\times 10^{-7}$ right after the initialization to $1.3351\times 10^{-9}$ at the last iteration, which indicates that VQE suffers from BPs heavily, hence can hardly achieve the ground energy. The dark-blue boxes in Fig. 8 and the dark-blue line in Fig. 9 plot these numerical results, which also match the results reported in Ref. [39].
@@
It turns out that right after the initialization, the variances of these two sets of gradients are $1.0695\times 10^{-2}$ and $1.2742\times 10^{-2}$, respectively, which are both five orders of magnitude larger than those in the original VQE method. When the optimizations are terminated, these variances eventually decrease to $7.8099\times 10^{-14}$ and $6.0972\times 10^{-6}$, respectively, with only VGON exhibiting a much larger variance magnitude than VQE. Fig. 8 illustrates the distribution of $\{|\partial_{{\theta_{l}^{i}}}h(\bm{\theta})|\}$ for the VQE-SA method with red boxplots, and that of $\{|\partial_{b_{l}^{i}}C(\bm{\phi,\omega})|\}$ for VGON with purple boxplots. As we can see, the absolute values of the gradients for the VQE-SA method and VGON are distributed more widely than those in VQE. Furthermore, a considerable part of these absolute values of gradients, especially at the initial stages, is several orders of magnitude larger compared to those in VQE, which is crucial for effectively decreasing the energies.
Figure 8: The absolute values of the gradients for the $Z_{1}Z_{2}$ model. The boxplots illustrate the distribution of $\{|\partial_{{\theta_{l}^{i}}}h(\bm{\theta})|\}$ for VQE (dark-blue) and the VQE-SA method (red), and the distribution of $\{|\partial_{b_{l}^{i}}C(\bm{\phi,\omega})|\}$ for the VGON (purple) at different iterations. Each boxplot displays the distribution based on a five-number summary: the minimum, the first quartile, the sampled median, the third quartile and the maximum. All other observed data points outside the minimum and maximum are plotted as outliers with black diamonds.
In addition, Fig. 9 illustrates the energies at different iterations for the two methods. As depicted by the red dashed line, the VQE-SA method converges to -1 fast, which is exactly the ground energy. In VGON, the average (minimal) energy is represented by the green (purple) dashed line, which also decreases rapidly, and eventually achieves a minimum value of -0.9998 (-0.9999). Compared with the VQE-SA method, at the beginning the fluctuations in VGON are smaller, which means VGON converges to the ground energy faster in this stage. However, since VGON tries to map the uniformly random parameters to those centered around the optimal parameters, there remain weak fluctuations in the later stage, but it still manages to find the ground energy very well.
Figure 9: Energies of the $Z_{1}Z_{2}$ model at different iterations for different methods. The VQE (dark-blue) suffers from BPs and can hardly be optimized. The VQE-SA method (red) and VGON, whose average (green) and minimal (purple) energies in each batch are presented, converge to the ground energy quickly. The exact ground energy (gray) is -1.
C.2 The Heisenberg XXZ model
@@
depicted in Fig. 10 (a) is applied. Here $L$ and $N$ are the numbers of the layers and qubits involved in $U^{\prime}(\bm{\theta})$ respectively, and each $U_{l}^{i}(\bm{\theta}_{l}^{i})$ is a univeral 2-qubit gate at the $l$-th layer acting on qubits $i$ and $i+1$, which is determined by $\bm{\theta}_{l}^{i}$ containing 15 rotation angles, as illustated in Fig. 10 (b).
@@
Figure 10: Structure of the parameterized quantum circuit for the Heisenberg XXZ model. (a) Each dark-blue box represents a layer of the circuit consisting of $n-1$ universal 2-qubit gate blocks. (b) Each universal 2-qubit gate block is decomposed into 15 rotation gates and 3 CNOT gates.
@@
When it turns to the VQE-SA method, a $\bm{\theta}$ is initialized uniformly from $\{\theta_{l}^{i}\in[0,0.01)\}$ as the starting point, and then it is updated iteratively to search for the ground state. As for VGON, 8,000 $\bm{\theta}$’s are uniformly initialized from $\{\theta_{l}^{i}\in[0,2\pi)\}$ as the inputs of the model, which contains a 7-layer encoder network with sizes [8192, 4096, 2048, 1024, 512, 256, 128], a latent space with dimension 100, and a 7-layer decoder network with sizes [128, 256, 512, 1024, 2048, 4096, 8192]. Set the batch size to 8, and the coefficient of the KL divergence to $0.1$. We run both of the two methods for 1,000 iterations with all the other configurations kept the same as VQE. To make a fair comparison, we repeat the whole process 10 times. Fig. 3 in the main text illustrates the corresponding mean values and the 95% confidence intervals of the energy densities and the fidelities between the optimized state and the ground state at different iterations, where we can clearly see the faster and more stable convergence of VGON than the VQE-SA method.
Appendix D Identifying degenerate ground state space of quantum models
@@
[…]
@@
In this section, we consider the ansatz expressed by Eq. (10), hence $M$ equals $15(N-1)L$, where $N$ and $L$ are the number of qubits and layers in the circuit, respectively.
D.1 The Majumdar-Ghosh model
@@
Figure 11: The overlaps between the generated states of VGON and the ground states. Each boxplot shows the degree of dispersion and skewness for the overlap with one ground state. Table 3 lists the minimum, lower quartile, median, upper quartile, and maximum for each ground state.
Table 3: Details on the overlap boxes of VGON’s output on the basis of the degenerate space of $H_{MG}$.
@@
Among the 1000 generated states, 81.9% achieve the energy threshold of -23.90. To examine the diversity of the generated states, we analyze the overlaps between the states achieving the above threshold and all the ground states, as shown in Fig. 11. It can be clearly seen that the generated states exhibit significant diversity. To ensure that VGON’s outputs effectively involve all the dimensions of the ground space, we set the overlap threshold to 0.001 and then analyze all the generated states. The results indicate that 81.4% of the generated states meet both the energy and the overlap thresholds. Fig. 12 exhibits ten such generated states, which not only illustrates the remarkable diversity of the solutions provided by VGON, but also demonstrates VGON’s capability of identifying degenerate ground state spaces for quantum models effectively.
Figure 12: The overlaps between selected generated states and the degenerate space. Five bases of the degenerate space, represented by $|v_{1}\rangle-|v_{5}\rangle$, are determined through the exact diagonalization method. For each generated state $|\psi_{i}\rangle$, the corresponding bars with different colors represent the overlaps into different bases.
D.2 The 232 model
@@
where $\sigma_{x,y,z}^{i}$ represent the Pauli matrices at site $i$. Let the number of qubits $N$ be 11, the corresponding ground energy is $-20.7106$.
Consider $L=6$, resulting in the number of universal 2-qubit gate blocks being 60, and the batch size $S_{b}=50$. The configurations of the VGON model and the training procedure remain consistent with that in Appendix D.1, except for setting the learning rate to 0.0015 and terminating the training upon reaching an energy value of -20.6106.
After training, 1,000 quantum states are generated by the VGON model, with 78.7% demonstrating energies below $-20.6106$. As illustrated in Fig. 13(a), the overlap distribution on each basis state, denoted as $\{|u_{1}\rangle,|u_{2}\rangle\}$, showcases a bimodal pattern, precisely reflecting the degree of degeneracy. Moreover, the analysis of fidelities between pairs of generated states, depicted in Fig. 13(b), reveals values clustering around either 0 or 1. This indicates that the states generated by the VGON model are either identical or orthogonal to each other. Consequently, this affirms VGONs’ capability to directly generate a complete set of orthogonal bases for this task.
Figure 13: Distributions of the generated states for the 232-type model. (a) The overlap distributions for the basis states $|u_{1}\rangle$ and $|u_{2}\rangle$, whose population densities plots are both bimodal, consistent with the degree of degeneracy. The minimum, lower quartile, median, upper quartile, and maximum for each basis state are presented in the first part of Table 4. (b) The distribution of fidelities between pairs of generated states with energies below -20.6106. The population densities plot reveals a bimodal distribution, with pronounced peaks near 0 and 1. This pattern suggests that the states are predominantly either identical (fidelity close to 1) or orthogonal (fidelity close to 0) to each other. The statistical summary of the boxplot is shown in the second part of Table 4.
Table 4: Boxplot details on overlap distributions for the basis of the degenerate space of $H_{232}$, and the distributions of fidelities between pairs of generated states.

### v2

@@ VII Conclusions
Using the default settings of the programs implemented by MATLAB, all the global optimization algorithms are executed on a computer with an Intel i9-12900KS Core and a RAM of 128 GB for 24 hours. Meanwhile, an MLP model with a batch size of 10, consisting of a 7-layer network with each layer containing 90 neurons is also trained 10 times with the other configuration remaining the same as that of VGON. The numerical results are shown in Supplementary Table 1, where we can clearly see that the performance of VGON exceeds that of all the other methods.
Supplementary Table 1: Comparisons of VGONs with seven global optimization algorithms and MLP.
@@
MLP 0.0384 (mean) 3,000 iterations
VGON 0.0837 3,000 iterations
Alleviating the effect of barren plateaux in variational quantum algorithms
@@
is adopted. Here $\theta_{l}^{i}\in[0,2\pi)$ are variational angles, and all the $L\times N$ such angles combined is denoted as $\bm{\theta}$. The rotations $R_{l}^{i}(\theta_{l}^{i})=\exp(-\frac{\rm{i}}{2}\theta_{l}^{i}G_{l}^{i})$ have random directions given by $G_{l}^{i}\in\{\sigma_{x},\sigma_{y},\sigma_{z}\}$. $U_{ENT}$ is an entangling unitary operation consisting of two-qubit nearest-neighbor controlled-Z (CZ) gates with periodic boundary conditions. $L$ and $N$ correspond to the numbers of the layers and qubits of $U(\bm{\theta})$ respectively. The structure of $U(\bm{\theta})$ for the $Z_{1}Z_{2}$ model is schematically shown in Supplementary Figure 3.
Supplementary Fig. 3: Structure of the parameterized quantum circuit for the $Z_{1}Z_{2}$ model. Each dark-blue wireframe represents a layer of the circuit consisting of single-qubit rotations represented by light-blue boxes and an entangling unitary operation $U_{ENT}$ represented as a purple box, where entangling CZ gates are shown by lines. The measurements at the end are used to estimate the energy of the trial state.
Let the numbers of qubits and layers be $20$ and $400$ respectively. To approximate the ground states of the $Z_{1}Z_{2}$ model by VQE, the variable $\bm{\theta}\in\mathbb{R}^{8,000}$ is uniformly initialized from the parameter space, i.e., $\theta_{l}^{i}\in[0,2\pi)$, and then updated iteratively by the Adam optimizer [34] to minimize $h(\bm{\theta})$. After 300 iterations, the obtained energy is -0.0068, which is actually far from the ground energy -1. Particularly, the variance of $\{\partial_{{\theta_{l}^{i}}}h(\bm{\theta})\}$ decreases from $4.4999\times 10^{-7}$ right after the initialization to $1.3351\times 10^{-9}$ at the last iteration, which indicates that VQE suffers from BPs heavily, hence can hardly achieve the ground energy. The dark-blue boxes in Supplementary Figure 4 and the dark-blue line in Supplementary Figure 5 plot these numerical results, which also match the results reported in Ref. [46].
@@
It turns out that right after the initialization, the variances of these two sets of gradients are $1.0695\times 10^{-2}$ and $1.2742\times 10^{-2}$, respectively, which are both five orders of magnitude larger than those in the original VQE method. When the optimizations are terminated, these variances eventually decrease to $7.8099\times 10^{-14}$ and $6.0972\times 10^{-6}$, respectively, with only VGON exhibiting a much larger variance magnitude than VQE. Supplementary Figure 4 illustrates the distribution of $\{|\partial_{{\theta_{l}^{i}}}h(\bm{\theta})|\}$ for the VQE-SA method with red boxplots, and that of $\{|\partial_{b_{l}^{i}}C(\bm{\phi,\omega})|\}$ for VGON with purple boxplots. As we can see, the absolute values of the gradients for the VQE-SA method and VGON are distributed more widely than those in VQE. Furthermore, a considerable part of these absolute values of gradients, especially at the initial stages, is several orders of magnitude larger compared to those in VQE, which is crucial for effectively decreasing the energies.
Supplementary Fig. 4: The absolute values of the gradients for the $Z_{1}Z_{2}$ model. The boxplots illustrate the distribution of $\{|\partial_{{\theta_{l}^{i}}}h(\bm{\theta})|\}$ for VQE (dark-blue) and the VQE-SA method (red), and the distribution of $\{|\partial_{b_{l}^{i}}C(\bm{\phi,\omega})|\}$ for the VGON (purple) at different iterations. Each boxplot displays the distribution based on a five-number summary: the minimum, the first quartile, the sampled median, the third quartile and the maximum. All other observed data points outside the minimum and maximum are plotted as outliers with black diamonds.
In addition, Supplementary Figure 5 illustrates the energies at different iterations for the two methods. As depicted by the red dashed line, the VQE-SA method converges to -1 fast, which is exactly the ground energy. In VGON, the average (minimal) energy is represented by the green (purple) dashed line, which also decreases rapidly, and eventually achieves a minimum value of -0.9998 (-0.9999). Compared with the VQE-SA method, at the beginning the fluctuations in VGON are smaller, which means VGON converges to the ground energy faster in this stage. However, since VGON tries to map the uniformly random parameters to those centered around the optimal parameters, there remain weak fluctuations in the later stage, but it still manages to find the ground energy very well.
Supplementary Fig. 5: Energies of the $Z_{1}Z_{2}$ model at different iterations for different methods. The VQE (dark-blue) suffers from BPs and can hardly be optimized. The VQE-SA method (red) and VGON, whose average (green) and minimal (purple) energies in each batch are presented, converge to the ground energy quickly. The exact ground energy (gray) is -1.
The Heisenberg XXZ model
@@
depicted in Supplementary Figure 6 (a) is applied. Here $L$ and $N$ are the numbers of the layers and qubits involved in $U^{\prime}(\bm{\theta})$ respectively, and each $U_{l}^{i}(\bm{\theta}_{l}^{i})$ is a univeral 2-qubit gate at the $l$-th layer acting on qubits $i$ and $i+1$, which is determined by $\bm{\theta}_{l}^{i}$ containing 15 rotation angles, as illustated in Supplementary Figure 6 (b).
@@
[…]
Supplementary Fig. 6: Structure of the parameterized quantum circuit for the Heisenberg XXZ model. (a) Each dark-blue box represents a layer of the circuit consisting of n1n-1 universal 2-qubit gate blocks. (b) Each universal 2-qubit gate block is decomposed into 15 rotation gates and 3 CNOT gates.
@@
When it turns to the VQE-SA method, a $\bm{\theta}$ is initialized uniformly from $\{\theta_{l}^{i}\in[0,0.01)\}$ as the starting point, and then it is updated iteratively to search for the ground state. As for VGON, 8,000 $\bm{\theta}$’s are uniformly initialized from $\{\theta_{l}^{i}\in[0,2\pi)\}$ as the inputs of the model, which contains a 7-layer encoder network with sizes [8192, 4096, 2048, 1024, 512, 256, 128], a latent space with dimension 100, and a 7-layer decoder network with sizes [128, 256, 512, 1024, 2048, 4096, 8192]. Set the batch size to 8, and the coefficient of the KL divergence to $0.1$. We run both of the two methods for 1,000 iterations with all the other configurations kept the same as VQE. To make a fair comparison, we repeat the whole process 10 times. Figure 3 in the main text illustrates the corresponding mean values and the 95% confidence intervals of the energy densities and the fidelities between the optimized state and the ground state at different iterations, where we can clearly see the faster and more stable convergence of VGON than the VQE-SA method.
Identifying degenerate ground state space of quantum models
@@
where $\mathcal{C}_{S_{b}}^{2}$ represents the set of all 2-combinations pairs derived from the elements in $\{1,2,\cdots,S_{b}\}$, and $\|\cdot\|$ denotes the Euclidean norm. Eventually, the optimization objective is set as minimizing of a combination of $\bar{E}(\bm{\Theta})$ and $\bar{S}_{\mathcal{C}_{S_{b}}^{2}}(\bm{\Theta})$ according to a trade-off coefficient $\gamma$, i.e.,
@@
In this section, we consider the ansatz expressed by Eq. (11), hence $M$ equals 15(N1)L15(N-1)L, where $N$ and $L$ are the number of qubits and layers in the circuit, respectively.
The Majumdar-Ghosh model
@@
Supplementary Fig. 7: The overlaps between the generated states of VGON and the ground states. Each boxplot shows the degree of dispersion and skewness for the overlap with one ground state. Supplementary Table 2 lists the minimum, lower quartile, median, upper quartile, and maximum for each ground state.
Supplementary Table 2: Details on the overlap boxes of VGON’s output on the basis of the degenerate space of $H_{MG}$.
@@
Among the 1000 generated states, 81.9% achieve the energy threshold of -23.90. To examine the diversity of the generated states, we analyze the overlaps between the states achieving the above threshold and all the ground states, as shown in Supplementary Figure 7. It can be clearly seen that the generated states exhibit significant diversity. To ensure that VGON’s outputs effectively involve all the dimensions of the ground space, we set the overlap threshold to 0.001 and then analyze all the generated states. The results indicate that 81.4% of the generated states meet both the energy and the overlap thresholds. Supplementary Figure 8 exhibits ten such generated states, which not only illustrates the remarkable diversity of the solutions provided by VGON, but also demonstrates VGON’s capability of identifying degenerate ground state spaces for quantum models effectively.
Supplementary Fig. 8: The overlaps between selected generated states and the degenerate space. Five bases of the degenerate space, represented by $|v_{1}\rangle-|v_{5}\rangle$, are determined through the exact diagonalization method. For each generated state $|\psi_{i}\rangle$, the corresponding bars with different colors represent the overlaps into different bases.
The 232 model
@@
where $\sigma_{x,y,z}^{i}$ represent the Pauli matrices at site $i$. Let the number of qubits $N$ be 11, the corresponding ground energy is 20.7106-20.7106.
Consider $L=6$, resulting in the number of universal 2-qubit gate blocks being 60, and the batch size $S_{b}=50$. The configurations of the VGON model and the training procedure remain consistent with that in Appendix The Majumdar-Ghosh model, except for setting the learning rate to 0.0015 and terminating the training upon reaching an energy value of -20.6106.
After training, 1,000 quantum states are generated by the VGON model, with 78.7% demonstrating energies below 20.6106-20.6106. As illustrated in Supplementary Figure 9(a), the overlap distribution on each basis state, denoted as $\{|u_{1}\rangle,|u_{2}\rangle\}$, showcases a bimodal pattern, precisely reflecting the degree of degeneracy. Moreover, the analysis of fidelities between pairs of generated states, depicted in Supplementary Figure 9(b), reveals values clustering around either 0 or 1. This indicates that the states generated by the VGON model are either identical or orthogonal to each other. Consequently, this affirms VGONs’ capability to directly generate a complete set of orthogonal bases for this task.
Supplementary Fig. 9: Distributions of the generated states for the 232-type model. (a) The overlap distributions for the basis states $|u_{1}\rangle$ and $|u_{2}\rangle$, whose population densities plots are both bimodal, consistent with the degree of degeneracy. The minimum, lower quartile, median, upper quartile, and maximum for each basis state are presented in the first part of Supplementary Table 3. (b) The distribution of fidelities between pairs of generated states with energies below -20.6106. The population densities plot reveals a bimodal distribution, with pronounced peaks near 0 and 1. This pattern suggests that the states are predominantly either identical (fidelity close to 1) or orthogonal (fidelity close to 0) to each other. The statistical summary of the boxplot is shown in the second part of Supplementary Table 3.
Supplementary Table 3: Boxplot details on overlap distributions for the basis of the degenerate space of $H_{232}$, and the distributions of fidelities between pairs of generated states.
@@
Neural network settings for different tasks
For clarity and brevity, we summarize the neural network hyperparameters used across these tasks in Supplementary Table 4. It can be found that the choice of latent space dimension and KL coefficient should align with the problem scale and optimization landscape complexity. Moderate settings suffice for simpler tasks—such as LTI models or pure/mixed-state cases—and for models with smoother landscapes like the Z1Z2 model. In contrast, more complex systems, including the XXZ model with larger state spaces and the MG model prone to trapping in degenerate ground-state subspaces, demand higher latent dimensions and KL coefficient annealing.
Small latent dimensions suffice when the intrinsic dimensionality of optimal solutions is low, even for challenging problems like mixed states. However, as the solution space grows more complex or higher-dimensional, the latent space must expand accordingly.
Supplementary Table 4: Neural network settings for different tasks including encoder structure $E$, decoder $D$ batch size $S_{b}$, latent dimension $z$ and KL coefficients $\beta$.
Tasks $E$ $D$ $S_{b}$ $z$ $\beta$
LTI model [8,4] [4,8,16] 2 2 1
Pure state case [512,256,128] [128,256,512] 6 2 1
Mixed state case [1024,512,256,128] [128,256,512] 5 2 1
$Z_{1}Z_{2}$model [256,128,64,32] [32,64,128,256] 4 3 1/8
XXZ model [8192,4096,…,256,128] [128,256,…,4096,8192] 8 100 0.1
MG model [512,256,128,64] [64,128,256,512] 50 50 1
232 model [512,256,128,64] [64,128,256,512] 50 50 1